# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

### v2

@@ Appendix H Hyperparameters
Appendix I Level Evolution
I.1 Visualization of Level Evolution in Minigrid
The visualization of how the Minigrid environment evolves as the number of blocks increases is shown in Figure 15. Each step of the evolutionary process produces an edited level that has a high learning efficiency.
I.2 Evolution of Levels in BipedalWalker
Figure 16 shows the complexity metric results trained by 3 methods, TRACED (Ours), ACCEL + ATPL, ACCEL. Starting with plain terrain (near zero point), all 3 methods guide levels with more complex terrain. With aspect of (a) Stump height, (b) Stump height high, (e) Stair height step metrics, the results show TRACED and ACCEL + ATPL quickly evolve levels compared to ACCEL. This rapid level increasing brings, short wall-clock relative to performance and high performance results in various test environments.
(a) Stump height low
(b) Stump height high
(c) Stair height low
(d) Stair height high
(e) Stair height step
(f) Pit gap low
(g) Pit gap high
(h) Roughness
Figure 16: Emergent BipedalWalker terrain complexity metrics. Aspect of (a) Stump height low, (b) Stump height high, and (e) Stair height step, complexity grows faster under TRACED (Ours, pink) or ACCEL + ATPL (brown) than under ACCEL (blue). The result is averaged over 5 random seeds. This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
I.3 Visualization of Level Evolution in BipedalWalker
Figure 17 shows the visualization of how the environment evolves as the complexity of the task increases. As we can see, the process starts from a plain terrain and gradually evolves to a level of terrain with increasingly complex parameters.
Appendix J Agent Trajectory Visualizations Across Environments
J.1 MiniGrid
Figure 18 shows a representative rollout in the PerfectMaze environment.
J.2 BipedalWalker
Figure 19 illustrates a zero-shot rollout in BipedalWalkerHardcore. The agent dynamically modulates joint torques to ascend and descend steps, leap over gaps, step onto stumps, and maintain balance on uneven ground, confirming TRACED’s effectiveness in teaching transferrable locomotion skills.
Appendix K Difference with Former Studies
K.1 Difference with ACCEL
K.2 Difference with CENIE
CENIE introduces a complementary curriculum strategy by explicitly measuring environmental novelty, rather than relying solely on regret signals as in PLR or ACCEL. It fits a Gaussian mixture model (GMM) to the agent’s past trajectories and scores new tasks by how poorly the GMM explains them, thereby steering the agent toward unfamiliar regions of the state–action space. This model-agnostic approach can be applied on top of any UED framework, including PLR and ACCEL.
Appendix L Related Work
Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments [24, 30, 4, 19]. PAIRED [4] trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments. However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula. To mitigate these limitations, Mediratta et al. [9] introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr [31] employs variational autoencoder–based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
Another line of work is based on Prioritized Level Replay (PLR) [5], which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks. This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent. PLR  [6] make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer. This leads to better performance on unseen environments counterintuitively. ACCEL [7] extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer. This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks. However, these methods [5, 6, 7] approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
Moving beyond regret-based approaches, CENIE [20] integrates novelty with regret to expose agents to more diverse learning situations. Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer. This curriculum-awareness leads to better exploration and generalization. DIPLR [32] introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state–action trajectories. MBeDED [33] introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task. It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task. So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains. For Multi-Agent Reinforcement Learning, MAESTRO [34] extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations. In contrast, RACCOON [10] applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration. For meta-RL, GROOVE [35] suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training. DIVA [36] is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage. In sim-to-real settings, GCL [37] aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task [38]. Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress [39]. Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process [40, 41]. WAKER [42] leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios. However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
Appendix M More Details On Baseline Algorithms
Robust PLR (PLR[6] was introduced as an extension of the original PLR [5]. While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay. When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks. The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer. Policy updates are exclusively performed on tasks sampled from the replay buffer. By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
Adversarially Compounding Complexity by Editing Levels (ACCEL) [7] actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space. This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret. ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation. The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities. The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid. This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity. Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
Coverage-based Evaluation of Novelty In Environment (CENIE) [20] introduces environment novelty as a complementary objective to regret in UED. CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum. The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space. At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences. Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution. The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state–action pairs and the learned distribution of past state–action experiences. CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) [19] frames environment generation as a diffusion process steered by the agent’s regret. Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase [29], and assessed ADD with the ADD codebase [19]. CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
Appendix N Long-term Analysis on TRACED
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 20: TRACED Solved-Rate Time Series. Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed. Each curve shows the instantaneous solved rate at each update step.
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 21: ACCEL Solved-Rate Time Series. Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed. Each curve shows the instantaneous solved rate at each update step.
TRACED reaches near-peak solved rates very quickly. However, its performance subsequently oscillates rather than remaining perfectly stable (Figure 20). Crucially, this same post-peak fluctuation also appears under the ACCEL curriculum (Figure 21), where the agent simply arrives more slowly at its maximum but then exhibits the identical up-and-down pattern. In other words, these oscillations reflect an inherent tendency of the RL policy to suffer occasional catastrophic failures in held-out levels, rather than a pathology introduced by our task-prioritization strategy. Addressing and damping these fluctuations in the student agent’s performance is an important direction for future work.
Appendix O Limitations
  • Co-Learnability In this paper, we proposed Co-Learnability, simple yet effective method. Follow-up studies could address these gaps by developing multi-step Co-Learnability estimators, incorporating importance-weighting or counterfactual techniques to reduce sampling bias, and combining observational co-learn signals with auxiliary models (e.g., task-conditioned value predictors) for stronger causal inference.
  • Sampling levels We sampled a fixed number of the lowest-regret levels as easy tasks to mirror ACCEL’s approach and enable fair comparisons with baseline methods. However, this simple strategy leaves room to explore more advanced task-selection mechanisms.
  • Heuristic weighting We introduced weighted approach to regret approximation and task prioritization that relies on two key parameters $\alpha$, PVL and ATPL in the regret approximation, and $\beta$, which weights the relative importance of Co-Learnability against task difficulty. While these parameters were manually tuned in the current implementation and demonstrated effective results across environments, future work could explore more principled methods for automatically determining these weights.
  • Level mutation We adopted the same approach as ACCEL which employs random mutations. Our approach focused on improving task priority rather than advancing environment parameter mutation mechanisms. The current random mutation process creates a clear opportunity for future work to advance these methods using evolutionary algorithms such as CMA-ES [43], so the Editor can offer the agent more informative tasks.
  • Extension to other RL algorithms Our experiments have been conducted with PPO-based student agents. While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum. Off-policy methods such as SAC [44] and TD3 [45] feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework. Moreover, Model-based approaches like DreamerV3 [46] build world models for planning, offering complementary insights into curriculum design. Evaluating these methods under our curriculum framework remains an important direction for future work.
  • Extend to other UED methods We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost. While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR [6] and CENIE [20]) and conduct comprehensive experiments to assess its performance in these settings.
  • Evaluation in other domains Our experimental validation has been conducted in MiniGrid [16, 4] and BipedalWalker [23, 7]. Evaluating TRACED in other domains such as Car Racing [23, 6], physical robotics applications [47], and reasoning task ARC [48] would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
Appendix P Experimental Setup and Reproducibility

### v3

@@ I.2 Difference with CENIE
Appendix P Numerical Results
P.1 Minigrid Environment
Table 11: MiniGrid Zero-Shot Solved Rates. Bold indicates the best result per task; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
20k 0.51 $\pm$ 0.13 0.86 $\pm$ 0.06 0.64 $\pm$ 0.12 0.65 $\pm$ 0.1
20k 0.27 $\pm$ 0.11 0.55 $\pm$ 0.15 0.18 $\pm$ 0.1 0.58 $\pm$ 0.14
20k 0.73 $\pm$ 0.02 0.8 $\pm$ 0.03 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02
20k 0.5 $\pm$ 0.02 0.53 $\pm$ 0.03 0.61 $\pm$ 0.03 0.49 $\pm$ 0.03
20k 0.74 $\pm$ 0.1 0.61 $\pm$ 0.11 0.74 $\pm$ 0.04 0.22 $\pm$ 0.13
20k 0.54 $\pm$ 0.13 0.32 $\pm$ 0.11 0.72 $\pm$ 0.06 0.25 $\pm$ 0.12
20k 0.58 $\pm$ 0.16 0.92 $\pm$ 0.06 0.9 $\pm$ 0.1 0.8 $\pm$ 0.13
20k 0.18 $\pm$ 0.12 0.64 $\pm$ 0.15 0.63 $\pm$ 0.14 0.78 $\pm$ 0.13
Maze 10k 0.02 $\pm$ 0.01 0.13 $\pm$ 0.1 0.04 $\pm$ 0.02 0.41 $\pm$ 0.13 0.7 $\pm$ 0.1
20k 0.1 $\pm$ 0.1 0.52 $\pm$ 0.16 0.2 $\pm$ 0.13 0.59 $\pm$ 0.14
20k 0.16 $\pm$ 0.11 0.56 $\pm$ 0.14 0.28 $\pm$ 0.1 0.65 $\pm$ 0.13
20k 0.46 $\pm$ 0.14 0.95 $\pm$ 0.04 0.54 $\pm$ 0.16 0.86 $\pm$ 0.09
20k 0.2 $\pm$ 0.05 0.55 $\pm$ 0.08 0.36 $\pm$ 0.08 0.53 $\pm$ 0.08
20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal et al. (2021). We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (each at 10k and 20k). At 10k, TRACED achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best in one maze (Maze2). Averaged across all twelve mazes, TRACED attains a mean solved rate of 0.66, outperforming all baseline configurations.
Table 12: MiniGrid Ablation Results. Bold indicates the best result per task; underline indicates the second-best.
Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
16Rooms 0.7 $\pm$ 0.19 0.72 $\pm$ 0.14 0.91 $\pm$ 0.06 0.79 $\pm$ 0.19
16Rooms2 0.44 $\pm$ 0.17 0.01 $\pm$ 0.01 0.79 $\pm$ 0.17 0.72 $\pm$ 0.17
SimpleCrossing 0.77 $\pm$ 0.02 0.68 $\pm$ 0.04 0.86 $\pm$ 0.04 0.89 $\pm$ 0.01
FourRooms 0.46 $\pm$ 0.04 0.36 $\pm$ 0.02 0.48 $\pm$ 0.04 0.47 $\pm$ 0.02
SmallCorridor 0.03 $\pm$ 0.03 0.44 $\pm$ 0.21 0.57 $\pm$ 0.2 0.49 $\pm$ 0.17
LargeCorridor 0.11 $\pm$ 0.1 0.42 $\pm$ 0.22 0.65 $\pm$ 0.21 0.5 $\pm$ 0.14
Labyrinth 0.55 $\pm$ 0.23 0.4 $\pm$ 0.24 0.49 $\pm$ 0.2 1.0 $\pm$ 0.0
Labyrinth2 0.88 $\pm$ 0.1 0.16 $\pm$ 0.11 0.62 $\pm$ 0.18 0.98 $\pm$ 0.01
Maze 0.42 $\pm$ 0.24 0.29 $\pm$ 0.19 0.63 $\pm$ 0.2 0.59 $\pm$ 0.17
Maze2 0.42 $\pm$ 0.24 0.3 $\pm$ 0.19 0.41 $\pm$ 0.15 0.42 $\pm$ 0.19
Maze3 0.8 $\pm$ 0.2 0.68 $\pm$ 0.2 0.86 $\pm$ 0.12 0.86 $\pm$ 0.07
PerfectMaze(M) 0.33 $\pm$ 0.04 0.34 $\pm$ 0.02 0.53 $\pm$ 0.07 0.66 $\pm$ 0.11
Mean 0.49 $\pm$ 0.04 0.4 $\pm$ 0.04 0.65 $\pm$ 0.03 0.7 $\pm$ 0.04
Table 12 presents zero-shot solved rates on twelve held-out MiniGrid mazes for TRACED and its ablations at 10k updates, averaged over 10 seeds (mean $\pm$ s.e.). Full TRACED achieves the highest score in six mazes (SimpleCrossing, Labyrinth, Labyrinth2, Maze3, PerfectMaze(M), and FourRooms) and ranks second in the remaining six, yielding a mean solved rate of 0.70 and outperforming every variant. These results indicate that both the transition-prediction error and Co-Learnability are critical to TRACED’s strong, consistent performance across diverse tasks.
Table 13: PerfectMaze Zero-Shot Solved Rates. Bold indicates the best result per environment; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
20k 0.02 $\pm$ 0.01 0.19 $\pm$ 0.09 0.18 $\pm$ 0.1 0.20 $\pm$ 0.08
20k 0.0 $\pm$ 0.0 0.12 $\pm$ 0.09 0.14 $\pm$ 0.09 0.09 $\pm$ 0.06
Table 13 compares zero‐shot solved rates of TRACED against DR, PLR, ADD, and ACCEL on two large‐scale PerfectMaze environments: PerfectMazeLarge and PerfectMazeXL, averaged over 10 seeds (mean $\pm$ s.e.). After 10k updates, TRACED achieves the highest success rate on PerfectMazeLarge (27% ± 7%) and the third‐highest on PerfectMazeXL (10% ± 4%), matching ADD at 20k. These results underscore TRACED’s ability to scale to extreme maze sizes.
P.2 BipedalWalker Environment
Table 14 compares TRACED’s zero-shot returns with DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains, averaged over five seeds (mean $\pm$ s.e.). ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Teoh et al. (2024) to our ACCEL implementation. At 10k updates, TRACED achieves the highest mean return on three terrains (Basic, Hardcore, and Stump) and attains an overall mean of $89.95\pm 31.72$, outperforming all baselines.
Table 14: BipedalWalker Zero-Shot Test Returns. Bold indicates the best result per terrain; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL ACCEL-CENIE TRACED
20k -1.55 $\pm$ 8.83 -1.89 $\pm$ 8.93 14.63 $\pm$ 21.99 62.69 $\pm$8.91 75.81
Table 15 reports zero-shot solved rates across six BipedalWalker terrains for DR, PLR, ADD, ACCEL, and TRACED, averaged over five seeds. A trajectory is considered “solved” if its return exceeds 230. At 10k updates, TRACED achieves the highest solved rate on five of the six tasks (Basic, Hardcore, Stairs, PitGap, Stump) and ranks second on Roughness, yielding an overall mean of 0.36, substantially higher than ACCEL’s 0.29 and the other baselines. No baseline matches TRACED’s combined efficiency and reliability in zero-shot generalization.
Table 15: BipedalWalker Zero‐Shot Solved Rates. Bold indicates the best per terrain; underline indicates the second‐best.
Environment Update DR PLR ADD ACCEL TRACED
20k 0.16 $\pm$ 0.13 0.35 $\pm$ 0.22 0.2 $\pm$ 0.2 0.99 $\pm$ 0.0
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.04 $\pm$ 0.04 0.23 $\pm$ 0.08
Stairs 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.02 0.03 $\pm$ 0.02
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.01
PitGap 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.02
Stump 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.18 $\pm$ 0.12
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.0.2 0.0 $\pm$ 0.0
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.03 0.76 $\pm$ 0.03
20k 0.03 $\pm$ 0.02 0.06 $\pm$ 0.04 0.05 $\pm$ 0.05 0.34 $\pm$ 0.03